OpenStudy (anonymous):

[8.02] B. Simplify: square root 100y+ square root 25y- square root 64y

OpenStudy (anonymous):

\[\sqrt{100y}+\sqrt{25y}-\sqrt{64y}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{100}=10 \sqrt{25}=5\sqrt{64}=8\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok i see, missing some commas, i got confused

OpenStudy (anonymous):

sorry, forgot them.

OpenStudy (anonymous):

\[\sqrt{100}=10\] \[ \sqrt{25}=5\] \[\sqrt{64}=8\]